The third method for graphing the solution set is to use the slope and y-intercept. Answer the following:

a) Find the slope and the y-intercept of the line formed by the following equation. Provide an in-between step: 9x - 3y = 15.

b) Explain how you would use the slope and y-intercept to graph the line represented by this equation.
1
Expert's answer
2020-02-28T10:25:07-0500

Given straight line is "9x-3y=15 ....(1)"

To find the slope and y intercept we reduce the equation in (1) to"y= mx+c"

From (1)

"-3y=15-9x"

Dividing with (-3) we get


"y=-5+3x"

Therefore, slope "m=3"

y intercept "c= -5"

b) To draw the graph of the straight line in (1), we need two points.

One point is "(0,-5)"

(Since the straight line cuts the y axis at "c=-5"

To find the second point we use slope "m=3"

Slope

"m=\>(rise)/(run)"

"= \>3 /1"

Therefore, rising 3 units above "-5," run will be "1"

"\Rightarrow" we get a point "(1,-2)".

Joining the points "(0,-5)" and "(1,-2)" we get the graph of the straight line "9x-3y=15."